OpenStudy (anonymous):

How do you solve for 2sinxcosx=2sinx on the interval [0,2pi]

OpenStudy (anonymous):

First divide both sides by \(2\sin x\), and you get: \[ \cos x =1 \]This is your first equation. Now, we can only do this when \[ 2\sin x\neq 0 \]Now we we plug \(2\sin x=0\) into the equation we get: \[ 0 \cos x = 0\implies 0=0 \] Which is a true statement so we need to solve \[ 2\sin x = 0 \]as well

OpenStudy (danjs):

Cos(theta) = 1 The cosine of an angle, is the X coordinate of the point (cos(t), sin(t)) on the unit circle. At what angles , is the X coordinate equal to 1 on the unit circle? Since the radius is 1, All of the radius is in the X direction, along the x axis. Zero Degrees

OpenStudy (danjs):

Same thought process for the second equation @wio mentioned above

OpenStudy (anonymous):

The 'proper way' is to subtract then factor: \[ 2\sin x\cos x = 2\sin x\\ 2\sin x \cos x - 2\sin x =0\\ 2\sin x(\cos x-1)=0 \]It leads to the same equations.

OpenStudy (anonymous):

Thank you @wio and @DanJS